A business discovers a linear relationship between the number of shirts it can sell and the price per shirt. In particular, 20,000 shirts can be sold at $19 each and 2000 of the same shirts can be sold at $55 each. Write the slope-intercept equation of the demand line through the ordered pairs (20,000 shirts, $19) and (2000 shirts, $55). Then determine the number of shirts that can be sold at $50 each.
Answer by stanbon(75887)

Write the slope-intercept equation of the demand line through the ordered pairs (20,000 shirts, $19) and (2000 shirts, $55). Then determine the number of shirts that can be sold at $50 each.
slope = (55-19)/(2000-20000)= 36/-18000= -1/500
Form of the linear is y=mx+b where y=55, x=2000, m=-1/500; find b.
55=-1/500 (2000) + b
55 = -4 +b
b=59
EQUATION: price of shirt =(-1/500)(number of shirts) +59
How many can be sold at $50 each?
50=(-1/500)x+59
-9=-(1/500)x
x=4500 shirts can be sold at $50 each.
Cheers,
Stan H.